OpenStudy (anonymous):

which point lies on the line defined by 3x+6y=2?

OpenStudy (anonymous):

are there points to chosen from?

OpenStudy (anonymous):

but first thing you'll have to do nonetheless is to find out y in a y=mx+b sort of format. so 3x+6y=2 isolate the y 6y=-3x+2 reduce y so that it could =1 y=(-3x+2)/6 y=-1/2x+1/3

OpenStudy (anonymous):

there is (0,2) (0,6) (1, -1/6) (1, -1/3)

OpenStudy (anonymous):

now that the equation has been solved, think you could figure it?

OpenStudy (anonymous):

i already knew how to put it into standard form, I just don't know how to get the actual answer and i have to show my work

OpenStudy (anonymous):

oh, I see, then um, well one way to do this is going through it one by one.... actually that's the only way I know. so I'd set x=0 and see what I get for y, if it happened to be 2 or 6 then that's my point respectively. if not then I'd go on and set x=1 and if the y value is one of those points then that be the point.

OpenStudy (anonymous):

@juan_9292 is absolutely right

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're welcome :)